# Clustering Document Parts: Detecting and Characterizing Influence Campaigns From Documents

### v1

@@ 5.3 Our approach
[…]
@@
We use S-BERT (Reimers and Gurevych, 2019) to embed document parts and then employ two clustering algorithms, i.e., KMEANS (MacQueen, 1967) and HDBSCAN Campello et al. (2015), to cluster the embedded document parts. HDBSCAN is a hierarchical extension of DBSCAN (Ester et al., 1996) with various optimization methods implemented Campello et al. (2015). Due to the curse of dimensionality, we find that HDBSCAN does not easily produce clusters without an embedding reduction algorithm in place. We use the state-of-the-art UMAP algorithm (McInnes et al., 2018) for this purpose.
@@
To contrast the baselines with our approach, we use the same two classification algorithms (FNN and XGBoost) to classify high-influence clusters. In addition to the 95 general linguistic features, there are 7 cluster-level features that are specific to our current pipeline: top-10 uni-gram text frequency, top-10 bi-gram text frequency, top-10 tri-gram text frequency, weighted n-grams text frequency, average cosine similarity, percentage of unique documents, and cluster size. Top-10 n-grams text frequency is the average ratio of texts containing the top-10 n-grams, whereas the weighted n-grams text frequency is the weighted sum of the aforementioned top-10 n-grams text frequencies, where the weights are simply given as $\frac{n}{\sum_{n^{\prime}=1}^{3}n^{\prime}}$ for $n\in\{1,2,3\}$. Average cosine similarity (ACS) is the average cosine similarity of all unique text pairs in a cluster:
$ACS=\frac{\sum_{i}^{m}\sum_{j\neq i}^{m}\text{cos\_sim}(\text{text}_{i},\text{text}_{j})}{m(m-1)}$
These 5 features are designed as “hard” (ngrams) and “soft” (ACS) measurements of topical coherence of a cluster. Percentage of unique documents and cluster size are just basic attributes of a cluster.
@@
For a single clustering experiment, we simply classify any document, whose parts occur in at least one high-influence cluster, as a high-influence document. This low threshold is to ensure that short documents would not be excluded from being identified as high-influence documents, since they may only be segmented into one part and thus have no more than one association with high-influence clusters.
As mentioned, we propose the aggregation of high-influence clusters from multiple clustering experiments to expand the search for high-influence documents. However, as a result of aggregation, the chance of misidentifying a high-influence document based on a single association with a high-influence cluster increases, since false positives in high-influence clusters also accumulate with aggregation. To regulate the false positive rate, we set the minimum number of associations of a document with high-influence clusters to be 1/5 of the number of high-influence clusters after aggregation for the document to be considered as high-influence.
Cluster aggregation
We run a total of 135 clustering experiments by varying related parameters for the two clustering algorithms we use, detailed in the Appendix B. The resulting clusters are aggregated both for training high-influence cluster classifier as well as selecting high-influence documents from high-influence clusters.
@@ 6 Results
Direct-document 20.2±${}_{\pm 2.2}$ 18.9±${}_{\pm 14.9}$ 17.1±${}_{\pm 8.6}$ 77.3±${}_{\pm 9.3}$ 37.9±${}_{\pm 7.9}$ 50.7±${}_{\pm 9.1}$
Document-level (mean) 0.3±${}_{\pm 0.1}$ 0.7±${}_{\pm 0.1}$ 0.4±${}_{\pm 0.1}$ 90.7±${}_{\pm 2.5}$ 25.4±${}_{\pm 3.3}$ 38.2±${}_{\pm 4.6}$
+ Aggregation $0.0{}_{\pm 0.0}$±${}_{\pm 0.0}$ ${}_{\pm 0.0}$ 94.1±${}_{\pm 0.7}$ 28.6±${}_{\pm 3.1}$ 43.8±${}_{\pm 3.8}$
Sentence-level (mean) 28.3±${}_{\pm 4.1}$ 44.1±${}_{\pm 4.7}$ 32.8±${}_{\pm 4.2}$ 69.4±${}_{\pm 10.9}$ 50.4±${}_{\pm 2.7}$ 56.7±${}_{\pm 4.1}$
+ Aggregation 74.5±${}_{\pm 16.4}$ 43.2±${}_{\pm 4.1}$ 54.3±${}_{\pm 7.7}$ 86.5±${}_{\pm 1.8}$ 70.7±${}_{\pm 2.4}$ 77.8±${}_{\pm 2.0}$
[…]
+ Aggregation 72.5±${}_{\pm 4.5}$ 40.0±${}_{\pm 5.7}$ 51.5±${}_{\pm 5.7}$ 81.1±${}_{\pm 3.5}$ 71.1±${}_{\pm 7.3}$ 75.5±${}_{\pm 3.8}$
[…]
+ Aggregation ${}_{\pm 4.6}$ 61.8±${}_{\pm 8.6}$ 63.1±${}_{\pm 6.0}$ 80.2±${}_{\pm 3.5}$ 71.4±${}_{\pm 1.8}$ 75.5±${}_{\pm 0.9}$
@@ 6.1 Main findings
[…]
@@
[…]
@@ 6.2 Error analysis
Table 4 reports the average number of errors the three best models, from the two baseline approaches and our approach, make on the test set for the five runs. The error counts are broken down according to media types.
@@
Clearly, clustering document parts helps overcome the limitations faced by the two baseline approaches, since the model recognizes influence campaigns in documents irrespective of the genre and have less than half false negatives, compared to the other two approaches. We identify 19 documents (15 FNs and 4 FPs) where the models across the five runs consistently misclassify. By using the keyword “bio”, we identify 3 of the 19 documents that may be mislabelled. For the rest 16 misclassifed documents, we hypothesize that the errors are mainly caused by two reasons. First, sentences do not necessarily reflect the theme of the document, which, for example, may make our model confuse documents exposing an influence campaign with one that spreads it. Second, none of the other techniques (e.g., SBERT, clustering) used in our pipeline are free of errors, which can propagate and ultimately lead to a wrong classification decision. We wish to improve our pipeline along these two directions in the future.
@@ 6.3 Threshold for models with aggregation
Nevertheless, as shown in Figs 1 and 2 from Appendix C, the classification F1 with aggregation is almost always better than without aggregation for most models. Unsurprisingly, the performance curve shows an upside down U-shape, a trade-off between precision and recall as we vary $\beta$. Setting $\beta=0.2$ is a conservative choice, which does not lead to the optimal performance. In the future, we would like to explore an automatic way of finding the optimal value for $\beta$.
@@ Acknowledgement
This material is based on work supported by the Defense Advanced Research Projects Agency (DARPA) under Contracts No. HR01121C0186, No. HR001120C0037, and PR No. HR0011154158. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of DARPA. Rambow gratefully acknowledges support from the Institute for Advanced Computational Science at Stony Brook University.
@@ Appendix A Data
Twitter 2,795 (2,098 + 697) 26.6±${}_{\pm 13.1}$
Forum 887 (744 + 143) 330.7±${}_{\pm 574.4}$
News 2,039 (1,735 + 304) 654.8±${}_{\pm 851.9}$
Blog 415 (340 + 75) 945.0±${}_{\pm 1,668.1}$
Reddit 371 (280 + 91) 69.3±${}_{\pm 131.7}$
Other 160 (137 + 23) 92.0±${}_{\pm 102.3}$
@@ Appendix C Results
Figs 1 and 2 show the performance variation of our models at different document part levels plus aggregation, as a function of the threshold $\beta$: the minimum number of times a document must be associated with a high-influence cluster in order to qualify as high-influence document, proportional to the total number of high-influence clusters available. The mean performance of these models on clusters from each clustering experiment is shown in dashed lines, as a baseline comparison.

### v2

@@ 5.3 Our approach
We break down a document into parts in three ways. (1) We use the sentences of the document as its parts (Sentence-level). We use the default sentence segmentation algorithm from spaCy V3.5.3. (Honnibal et al., 2020) (2) We also experiment with the state-of-the-art event factuality prediction system (Murzaku et al., 2023) to extract from each sentence of a document (source, target, factuality label) triplets. Here, source refers to the belief holder, target is a head word denoting an event, and factuality label describes the extent to which the source believes that the event has happened, is happening, or will happen. The source can either be the author herself, or somebody else according to the author. The factuality label has five possible values, ranging from committed belief (certain that true) to committed disbelief (certain that false), with possible belief, unknown belief, and possible disbelief in between. We use a head-to-span algorithm to extract a multi-word text span, of which the target is the syntactic head, as the representation of the identified event to be used as the extracted document parts. For TargetALL-level, we use all target spans extracted by the belief system. (3) We use the same event factuality prediction system but we retain only the events believed by the author (TargetAT-level), a subset of all the events identified by the event factuality prediction system, to see if document parts to which the author holds a belief will lead to a better result using our approach. The examples in Table 1 are events believed by the author, where the target words are highlighted in bold. The number of sentences and targets in the train and test sets is listed in Table 2.
@@
We use S-BERT (Reimers and Gurevych, 2019) to embed document parts and then employ two clustering algorithms, i.e., KMEANS (MacQueen, 1967) and HDBSCAN Campello et al. (2015), to cluster the embedded document parts. HDBSCAN is a hierarchical extension of DBSCAN (Ester et al., 1996) with various optimization methods implemented Campello et al. (2015). Due to the curse of dimensionality Bellman (1961), HDBSCAN does not easily produce clusters without an embedding reduction algorithm in place. We use the state-of-the-art UMAP algorithm (McInnes et al., 2018) for this purpose.
@@
To contrast the baselines with our approach, we use the same two classification algorithms (FNN and XGBoost) to classify high-influence clusters. In addition to the 95 general linguistic features, there are 7 cluster-level features that are specific to our current pipeline: top-10 uni-gram text frequency, top-10 bi-gram text frequency, top-10 tri-gram text frequency, weighted n-grams text frequency, average cosine similarity between all pairs of document parts in the cluster, percentage of unique documents, and cluster size. Top-10 n-grams text frequency is the average ratio of texts containing the top-10 n-grams, whereas the weighted n-grams text frequency is the weighted sum of the aforementioned top-10 n-grams text frequencies, where the weights are simply given as $\frac{n}{\sum_{n^{\prime}=1}^{3}n^{\prime}}$ for $n\in\{1,2,3\}$. Average cosine similarity (ACS) is the average cosine similarity of all unique text pairs in a cluster:
$\mathrm{ACS}=\frac{\sum_{i}^{m}\sum_{j\neq i}^{m}\text{cos\_sim}(\text{text}_{i},\text{text}_{j})}{m(m-1)}$
These 5 features are designed as “hard” (ngrams) and “soft” (ACS) measurements of topical and thematic coherence of a cluster, which are relational and independent of the specific lexical choices used inside the cluster. Percentage of unique documents and cluster size are just basic attributes of a cluster. The percentage of unique documents is calculated by dividing the number of documents whose parts occur in the cluster by the number of document parts in the cluster.
Cluster aggregation
We run a total of 135 clustering experiments by varying related parameters for the two clustering algorithms we use, detailed in the Appendix B. We use all resulting clusters both for training the high-influence cluster classifier, as well as for selecting high-influence documents from high-influence clusters.
@@
For a single clustering experiment, we simply classify any document, whose parts occur in at least one high-influence cluster, as a high-influence document. This low threshold is to ensure that short documents would not be excluded from being identified as high-influence documents, since they may only be segmented into one part and cannot have more than one association with high-influence clusters.
As mentioned, we propose using all high-influence clusters from multiple clustering experiments to expand the search for high-influence documents. However, as a result of this aggregation, the chance of misidentifying a high-influence document based on a single association with a high-influence cluster increases, since false positives in high-influence clusters also accumulate with aggregation. To regulate the false positive rate, we set $\beta=0.2$ (see Sec 3 for definition).
Evaluation
Since a clustering configuration does not necessarily produce high-influence clusters and in practice we can always try to find one that does, we evaluate the average performance of our approach on various clustering setups that produce high-influence clusters predicted by our pipeline. We choose a wide range of clustering configurations (see Appendix B) so as to avoid hard finetuning our approach. We run the two classification algorithms (FNN and XGBoost) for five times with varying parameters to compare the average performance of our approach against the baseline approaches on our dataset.
@@ 6 Results
Direct-document 20.2±2.2 18.9±14.9 17.1±8.6 77.3±9.3 37.9±7.9 50.7±9.1
Document-level (mean) 0.3±0.1 0.7±0.1 0.4±0.1 90.7±2.5 25.4±3.3 38.2±4.6
+ Aggregation 0.0±0.0 0.0±0.0 0.0±0.0 94.1±0.7 28.6±3.1 43.8±3.8
Sentence-level (mean) 28.3±4.1 44.1±4.7 32.8±4.2 69.4±10.9 50.4±2.7 56.7±4.1
+ Aggregation 74.5±16.4 43.2±4.1 54.3±7.7 86.5±1.8 70.7±2.4 77.8±2.0
TargetALL-level (mean) 25.4±6.7 35.2±8.5 27.0±6.9 78.2±3.7 73.8±2.4 75.3±1.3
+ Aggregation 72.5±4.5 40.0±5.7 51.5±5.7 81.1±3.5 71.1±7.3 75.5±3.8
TargetAT-level (mean) 60.7±7.1 66.8±10.5 62.4±8.5 63.5±2.2 49.5±2.8 54.8±2.1
+ Aggregation 64.8±4.6 61.8±8.6 63.1±6.0 80.2±3.5 71.4±1.8 75.5±0.9
@@ 6.1 Main findings
There is value in clustering belief targets. They are multi-word text spans within a sentence that carry a factuality label and involve a belief source. Compared to full sentences, they are more information-dense. We find that when FNN is used to classify clusters, clustering belief targets the author holds to be true (TargetAT-level) leads to the best performance, independent of the use of aggregation. When XGBoost is used, clustering all belief targets outperforms clustering sentences by 18% absolute in F1 without aggregation. These results show the potential of extracting belief targets for a better detection of influence campaigns, which intuitively make sense because influence campaigns are all about spreading a certain belief of the influencers.
@@
From Table 3, we see that cluster aggregation helps in every experiment (leaving aside document-level clustering using FNN, which performs at near-0 levels). Most of these improvements are statistically significant, given the standard deviations shown. However, for FNN using TargetAT-level there is no significant difference, and for XGBoost using TargetALL-level there is no significant difference. We have no explanation for these exceptions for now. In general, cluster aggregation helps in two ways. When our models have very low precision (using FNN), the current aggregation setup rules out many false positives resulting from misclassified high-influence clusters, which greatly improves precision. On the other hand, when precision is decent (in the case of XGBoost), aggregation can serve to help increase the range of relevant documents associated with high-influence clusters, which lead to a better recall in most cases.
@@ 6.2 Error analysis
Table 4 reports the average number of errors made by the two baseline approaches and our approach on the test set, over the five runs. The error counts are broken down according to media types.
@@
Clearly, clustering document parts helps overcome the limitations faced by the two baseline approaches, since the model recognizes influence campaigns in documents irrespective of the genre and have less than half false negatives, compared to the other two approaches. We identify 19 documents (15 FNs and 4 FPs) where the models across the five runs consistently misclassify. By using the keyword “bio”, we identify 3 of the 19 documents that may be mislabelled. For the remaining 16 misclassifed documents, we hypothesize that the errors are mainly caused by two reasons. First, sentences do not necessarily reflect the theme of the document, which, for example, may make our model confuse documents exposing an influence campaign with one that spreads it. Second, none of the other techniques (e.g., SBERT, clustering) used in our pipeline are free of errors, which can propagate and ultimately lead to a wrong classification decision. We wish to improve our pipeline along these two directions in the future.
@@ 6.3 Threshold for models with aggregation
Nevertheless, as shown in Fig 1 (also see Fig 2 from Appendix C for FNN), the classification F1 with aggregation is almost always better than without aggregation for most models. Unsurprisingly, the performance curve shows an upside down U-shape, a trade-off between precision and recall as we vary $\beta$. Setting $\beta=0.2$ is a conservative choice, which does not lead to the optimal performance. In the future, we would like to explore an automatic way of finding the optimal value for $\beta$.
@@ 7 Conclusion
There are several avenues for possible future work and we list three below. (1) Datasets. Given the increasing importance of detecting influence campaigns, we hope there will be more datasets annotated on the document collection level for an influence campaign. (2) Incorporating non-textual information. Our current pipeline is a text-only system. Leveraging non-textual information, such as social interactions and the authors’ past activities, may help us create a more complicated and comprehensive system (e.g., using graph neural network) that enhances the accurate and reliable detection of influence campaigns. However, such work cannot be possible without good datasets. (3) Automatic characterization of influence campaigns. Our work captures influence campaigns by the high-influence clusters, which may contain a large number of semantically related document parts, possibly with noise. To fully make sense of these clusters, we need to have some automatic ways of characterizing them in a fine-grained and interpretable way aligned with the downstream needs. Our preliminary experiments show that LLMs may be a potential option.
@@ Acknowledgement
We thank three anonymous reviewers from the 6th NLP+CSS Workshop for the constructive and helpful comments. This material is based on work supported by the Defense Advanced Research Projects Agency (DARPA) under Contracts No. HR01121C0186, No. HR001120C0037, and PR No. HR0011154158. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of DARPA. Rambow gratefully acknowledges support from the Institute for Advanced Computational Science at Stony Brook University.
@@ Appendix A Data
Twitter 2,795 (2,098 + 697) 26.6±13.1
Forum 887 (744 + 143) 330.7±574.4
News 2,039 (1,735 + 304) 654.8±851.9
Blog 415 (340 + 75) 945.0±1,668.1
Reddit 371 (280 + 91) 69.3±131.7
Other 160 (137 + 23) 92.0±102.3
@@ Appendix C Results
Fig 2 shows the performance variation of our models at different document part levels plus aggregation, as a function of the threshold $\beta$: the minimum number of times a document must be associated with a high-influence cluster in order to qualify as high-influence document, proportional to the total number of high-influence clusters available. The mean performance of these models on clusters from each clustering experiment is shown in dashed lines, as a baseline comparison.